# An information metric for comparing and assessing informative interim decisions in sequential clinical trials

### v2

@@ 2.3.1 Posterior distributions with normal data
Let $(f_{1},\dots,f_{s})$ and $(e_{1},\dots,e_{s})$ be the vectors of futility and efficacy stopping boundaries for the first $s$ interim analyses and let $g_{(s)}$ be the density of $s$ cumulative means, namely a multivariate normal distribution centered in $\bm{\theta}=(\theta,\dots,\theta)^{\prime}$ with elements of the covariance matrix given by $Cov(\bar{X}_{(s_{1})},\bar{X}_{(s_{2})})=\sigma^{2}/n_{(s_{2})}$ for each $s_{1}\leq s_{2}$. Then, the posterior distribution conditional on all the informative decisions made up to the end of stage $s$ is given by (2), with
@@ 2.3.2 Example: 3-stage group-sequential trial
Figure 1 shows conditional and unconditional posterior distributions for several scenarios, with different possible decisions occurring at the end of each stage. Unconditional posterior distributions at each stage have the same variance, that is affected only by the sample size, not by the observed values. In contrast, conditional posteriors vary much across scenarios within the same time. For example, observing a sample mean equal to either $0.5$ or $1$ at interim 1 (bottom and middle left panel, respectively) only shifts the mode of the unconditional posterior, while the conditional one also reflects the interim decision. This decision increases the posterior uncertainty on $\theta$ in both conditional posterior distributions, though to different extents. Specifically, stopping for efficacy at interim 1 after observing $\bar{x}_{(1)}>e_{1}$ yields a conditional posterior with much higher uncertainty, since $\bar{x}_{(1)}$ is close to the boundary $e_{1}$.
@@ 4.1 Numerical illustration: pre-experimental assessment of decision boundaries
We consider a three-stage group-sequential trial to test a single dose of a drug to treat disorders of the central nervous system, which was described by Dmitrienko and Koch, (2017). The primary objective of the trial is to demonstrate that the chosen dose of the drug was superior to a placebo in the acute treatment of subjects who meet the criteria for major depression. The efficacy of the experimental drug is evaluated using the mean reduction from baseline to the end of the 8-week study period in the total score of the 17-item Hamilton Depression Rating Scale (HAMD17). The standardized treatment effect is considered the primary endpoint, assuming that it is normal with mean $\delta$ and known standard deviation $\sigma=1$. The two interim analyses are scheduled to occur after $50\%$ and $75\%$ of the patients have completed the study.
The original efficacy boundaries of Pocock, (1977) and O’Brien and Fleming, (1979) assume equally spaced interim analyses. To accomodate for unequal spacing, we consider Pocock-type and O’Brien-Fleming-type efficacy boundaries implemented via their corresponding $\alpha$-spending functions (Lan and DeMets,, 1983) to maintain $\alpha\leq 0.025$ under $H_{0}:\delta\leq 0$ and power $1-\beta=0.9$ under $H_{1}:\delta=0.265$. Since the original analysis was not Bayesian, we consider a weakly-informative prior specification for $\delta\sim N(0,5)$, whose variance is $5$ times lower than one patient’s response. Figure 5 shows the expected sample size and the expected end-of-study AIPD associated with these two efficacy boundary options as a function of the unknown parameter $\delta$. O’Brien-Fleming-type boundaries are associated with both smaller expected sample size and expected AIPD for $\delta\leq 0.2$, which suggests a preference for these stopping boundaries over the Pocock-type boundaries when small treatment effects are anticipated by the clinicians. On the other hand, a sweet spot for Pocock-type stopping boundaries is represented by $\delta\in[0.25,0.4]$, although the gain in terms of expected number of patients recruited is limited. The expected AIPD is maximum around $\delta=0.225$ with Pocock-type boundaries and $\delta=0.325$ with O’Brien-Fleming-type boundaries. Intuitively, values of $\delta$ close to the first efficacy boundary $e_{1}$ are more likely to produce values of $\bar{x}_{(1)}$ that are close to this stopping boundary, thus inducing a higher AIPD due to the higher uncertainty about the decision made (see Figure 2). Unless clinicians are particularly concerned about a potential high conflict between conditional and unconditional posterior estimates - possibly leading to conditional posterior credible intervals that do not align with the final decision made in certain regions of the parameter space -, the smaller area under $\bar{D}(\theta)$ for Pocock-type boundaries compared to O’Brien-Fleming-type boundaries (13.9 vs. 16.3) suggests a lower expected divergence between the two posterior distributions when using the former set of boundaries.
@@
Figure 5: Expected sample size (left panel) and expected end-of-study adaptation-induced posterior divergence, $\bar{D}(\theta)$, (right panel) when using Pocock-type and O’Brien-Fleming-type efficacy boundaries in a 3-stage trial with normal endpoint. Stage-wise boundaries are $e_{P}=(0.25,0.20,0.17)$ and $e_{BF}=(0.33,0.22,0.16)$. Both boundary options guarantees a maximum type-I error probability $\alpha=0.025$ under $H_{0}:\delta\leq 0$ (solid line) and power $1-\beta=0.9$ under $H_{1}:\delta=0.265$ (dashed line). Pocock-type and O’Brien-Fleming-type are associated with a maximum sample size of 167 and 153 patients, respectively.
@@ 5 Discussion
We have considered cases where stopping rules are informative; thus, ignoring whether interim data meet or do not meet stopping criteria would neglect the informativeness of the stopping process (Roberts,, 1967; Berger and Wolpert,, 1988). This highlights the key distinction between conditional-on-decisions and unconditional posteriors. The conditional posterior treats informative stopping decisions as part of the data. In contrast, the unconditional (or counterfactual) posterior considers only the observed responses, treating interim decisions as non-informative. This should not be confused with the frequentist “unconditional” inference of Berry, (1987), which averages over all possible experiments: consistent with the Bayesian framework of Flournoy and Tarima, (2023), we always condition on the events that actually occurred.
@@
The expected version of the AIPD, $\bar{D}(\theta)$, can be computed before a trial starts to assess whether a given set of decision boundaries is likely to lead to large distortions in inference. This measure varies with $\theta$ and, thus, it can be reported in a similar way as the expected sample size. Additionally, the area below the curve of $\bar{D}(\theta)$ gives a global summary of the expected AIPD across all the possible scenarios. This measure can guide trialists on the choice of decision boundaries on the basis of the impact that interim decisions are anticipated to have on posterior estimates.

### v3

@@ 2.3.1 Posterior distributions with normal data
Let $(f_{1},\dots,f_{s})$ and $(e_{1},\dots,e_{s})$ be the vectors of futility and efficacy stopping boundaries for the first $s$ interim analyses and let $g_{(s)}$ be the density of $s$ cumulative means, namely a multivariate normal distribution centered in $\bm{\theta}=(\theta,\dots,\theta)^{\prime}$ with elements of the covariance matrix given by $Cov(\bar{X}_{(s_{1})},\bar{X}_{(s_{2})})=\sigma^{2}/n_{(s_{2})}$ for each $s_{1}\leq s_{2}$. Then, the posterior distribution conditional on all the informative decisions made up to the end of stage $s$ is given by (2), with (1) being
@@ 2.3.2 Example: 3-stage group-sequential trial
Figure 1 shows conditional and unconditional posterior distributions for several scenarios, with different possible decisions occurring at the end of each stage. Unconditional posterior distributions at each stage have the same variance, that is affected only by the sample size, not by the observed summary statistics. In contrast, conditional posteriors vary much across scenarios within the same time. For example, observing a sample mean equal to either $0.5$ or $1$ at interim 1 (bottom and middle left panel, respectively) only shifts the mode of the unconditional posterior, while the conditional one also reflects the interim decision. This decision increases the posterior uncertainty on $\theta$ in both conditional posterior distributions, though to different extents. Specifically, stopping for efficacy at interim 1 after observing $\bar{x}_{(1)}>e_{1}$ yields a conditional posterior with much higher uncertainty, since $\bar{x}_{(1)}$ is close to the boundary $e_{1}$.
@@ 4.1 Numerical illustration: pre-experimental assessment of decision boundaries
We consider a three-stage group-sequential trial to test a single dose of a drug to treat disorders of the central nervous system, which was described by Dmitrienko and Koch, (2017). The primary objective of the trial is to demonstrate that the chosen dose of the drug was superior to a placebo in the acute treatment of subjects who meet the criteria for major depression. The efficacy of the experimental drug is evaluated using the mean reduction from baseline to the end of the 8-week study period in the total score of the 17-item Hamilton Depression Rating Scale (HAMD17). The standardized treatment effect is considered the primary endpoint, assuming that it is normal with mean $\delta$ and known standard deviation $\sigma=1$. Since the original analysis was not Bayesian, we consider a weakly-informative prior specification for $\delta\sim N(0,5)$, whose variance is $5$ times greater than one patient’s response.
Figure 5 shows the expected sample size and the expected end-of-study AIPD associated with Pocock-type and O’Brien-Fleming-type efficacy boundaries under three alternative schedules of interim analyses. For example, under the first schedule, the two interim analyses are conducted after roughly $25\%$ and $50\%$ of patients have been recruited, respectively. Pocock-type and O’Brien-Fleming-type efficacy boundaries are implemented via their corresponding $\alpha$-spending functions (Lan and DeMets,, 1983) to accommodate for unequal spacing between interim analyses while maintaining $\alpha\leq 0.025$ under $H_{0}:\delta\leq 0$ and power $1-\beta=0.9$ under $H_{1}:\delta=0.265$. In all scenarios, O’Brien-Fleming-type boundaries are associated with both smaller expected sample size and smaller expected AIPD for small treatment effects ($\delta\leq 0.2$), making them preferable when modest efficacy is anticipated. When the true treatment effect is large ($\delta>0.5$), Pocock-type boundaries lead to smaller expected sample size and expected AIPD, with the sweet spot being most evident when the first interim analysis is scheduled early (e.g., at $25\%$ of information). When the first interim analysis is delayed (e.g., $50\%$ of information), differences between the two boundary types are less sharp, with expected AIPD being negligible when $\delta>0.5$. Notably, the expected AIPD under O’Brien-Fleming-type boundaries is more sensitive to the interim schedule: for instance, earlier first looks tend to shift the peak of the expected AIPD towards larger values of $\delta$. The area under $\bar{D}(\theta)$, which averages performance across all possible $\delta$, is consistently smaller for Pocock-type boundaries (ranging between 13.8 and 19.5, compared with 15.9 to 22.2 for O’Brien-Fleming-type), suggesting that they reduce the expected divergence between conditional and unconditional posteriors. Overall, O’Brien-Fleming-type seems to be preferable when only modest treatment effects are expected, while Pocock-type boundaries are recommended when larger effects are more plausible or when global performance across $\delta$ is prioritised.
@@
Figure 5: Expected sample size (top panels) and expected end-of-study adaptation-induced posterior divergence (bottom panels), $\bar{D}(\theta)$, when using Pocock-type and O’Brien-Fleming-type efficacy boundaries in a 3-stage trial with normal endpoint. Each column considers different schedules of interim analyses. All boundary options guarantee a maximum type-I error probability $\alpha=0.025$ under $H_{0}:\delta\leq 0$ (solid line) and power $1-\beta=0.9$ under $H_{1}:\delta=0.265$ (dashed line). Maximum sample size ranges from 167 to 176 for Pocock-type and 151 to 153 for O’Brien-Fleming-type boundaries across the three interim schedules.
@@ 5 Discussion
We have considered cases where stopping rules are informative; thus, ignoring whether interim data meet or do not meet stopping criteria would neglect the informativeness of the stopping process (Roberts,, 1967; Berger and Wolpert,, 1988, p. 88). This highlights the key distinction between conditional-on-decisions and unconditional posteriors. The conditional posterior treats informative stopping decisions as part of the data. In contrast, the unconditional (or counterfactual) posterior considers only the observed responses, treating interim decisions as non-informative. This should not be confused with the frequentist “unconditional” inference of Berry, (1987), which averages over all possible experiments: consistent with the Bayesian framework of Flournoy and Tarima, (2023), we always condition on the events that actually occurred.
@@
The expected version of the AIPD, $\bar{D}(\theta)$, can be computed before a trial starts to assess whether a given design configuration is likely to lead to large distortions in inference. This measure varies with $\theta$ and, thus, it can be reported in a similar way as the expected sample size. Additionally, the area below the curve of $\bar{D}(\theta)$ gives a global summary of the expected AIPD across all the possible scenarios. This measure can help trialists in selecting alternative combinations of decision boundaries and interim schedules, based on the anticipated impact of interim decisions on posterior estimates.